OpenStudy (anonymous):

4 + 4 * log(2) x = 14 with 2 as the base of the log

OpenStudy (anonymous):

t says to rund ur answer to 2 decimal places

OpenStudy (anonymous):

the answer was 5.66

OpenStudy (anonymous):

4logx=10 log x^4=10 x^4=2^10 x^2=2^5=32 x=sqrt(32)=5.66